# Homeomorphism groups of basilica, rabbit and airplane Julia sets

### v1

@@ 4.2 Embedding trees into dendrites
For all $n\geq 3$, there exist color-preserving topological embeddings $\Phi\colon\overline{T_{n}}\to D_{n}$ such that the closure of $\bigcup_{\{v,w\}\in E(T_{n})}\overline{C_{\Phi(v),\Phi(w)}}$ is $D_{n}$.
@@
Let $\Phi_{*}\colon V(T_{n})\to\mathrm{Br}(D_{n})$ be defined as $\Phi_{m}(v)$ for $m$ such that $v\in B_{v_{0}}(m)$. For each edge $\{v,w\}$ of $T_{n}$, extend $\Phi_{*}$ on the corresponding arc $[v,w]$ of $\overline{T_{n}}$ to be any homeomorphism from $[v,w]$ to the unique arc $[\Phi_{*}(v),\Phi_{*}(w)]$ of $D_{n}$. Thanks to condition (1) on each $\Phi_{m}$, this defines an embedding $\overline{T_{n}}\setminus\partial(T_{n})\to D_{n}\setminus\partial D_{n}$, which thus extends to an embedding $\Phi\colon\overline{T_{n}}\to D_{n}$. Thanks to condition (2) on the maps $\Phi_{m}$, the embedding $\Phi$ is color-preserving in the sense of Definition 4.7.
@@ 4.3 Embedding universal groups into kaleidoscopic groups
Given $g\in\mathcal{U}(N)$, we want to build $\varphi(g)\in\mathcal{K}(N)$. We start by considering $f\coloneqq\Phi\circ g\circ\Phi^{-1}$ on $\Phi(V(T_{n}))$. Note that $f$ is a partial dendrite homomorphism (Definition 4.2) when restricted to any finite subset of $\Phi(V(T_{n}))$. We will now extend $f$ “canonically” everywhere else using Lemma 4.4, as explained below.
Consider the set $\Omega_{\Phi(V(T_{n}))}$ of connected components of $D_{n}\setminus\Phi(V(T_{n}))$, which we will denote by $\Omega$ for the sake of brevity. Recall that the boundary of each element of $\Omega$ in $D_{n}$ consists of either one or two elements of $\Phi(V(T_{n}))$ by Remark 4.3. Actually, since $\Phi$ is such that the closure of $\bigcup_{\{v,w\}\in E(T_{n})}\overline{C_{\Phi(v),\Phi(w)}}$ is $D_{n}$ (Lemma 4.8), the boundary of each element $A$ of $\Omega$ must have two points, otherwise $A$ would be left out of such set. For the sake of brevity, given an edge $\{v,w\}$ of $T_{n}$, let us write $C_{v,w}$ for $C_{\Phi(v),\Phi(w)}$, which is the element of $\Omega$ whose boundary is $\{\Phi(v),\Phi(w)\}$. Let $D_{v,w}$ denote the closure of $C_{v,w}$, which is $C_{v,w}\cup\{\Phi(v),\Phi(w)\}$. Since $D_{v,w}$ is homeomorphic to $D_{n}$, for all $\{v,w\}\in E(T_{n})$ we can apply Proposition 4.6 to find color-preserving homeomorphisms $\psi_{(v,w)}\colon D_{v,w}\to D_{n}$ such that
@@
Since $\Phi$ is color-preserving (Definition 4.7) and $c_{b_{0}}|_{\Phi(HT(T_{n}))}$ is a bijection $\Phi(HT(T_{n}))\to\widehat{b_{0}}$, we have
$k_{T}=k_{D}\circ c_{b_{0}}\circ\Phi\circ\tau,$
@@
It is likely that statements similar to Lemmas 4.8, LABEL:, 4.9, LABEL: and 4.11 also hold for universal groups $\mathcal{U}(N,M)$ on biregular trees $T_{n,m}$ and generalized kaleidoscopic groups $\mathcal{K}(N,M)$ on Ważewski dendrites $D_{n,m}$. However, the groups $\mathcal{K}(N,M)$ have not been formally defined in the literature. They can be built by using two disjoint sets of colors $[n]$ and $[m]$ and fixing two local groups $N$ and $M$ acting on $[n]$ and $[m]$, respectively, as noted in [DMW19, Remark 3.10].
@@ 4.4 Embeddings and non-embeddings among the homeomorphism groups
In the case of dendrites and kaleidoscopic groups, let us see that again $\chi$ induces color-preserving embeddings $\mathcal{X}\colon D_{n_{1}}\hookrightarrow D_{n_{2}}$. We will first produce an injective map $\tau\colon\mathrm{Br}(D_{n_{1}})\to\mathrm{Br}(D_{n_{2}})$ that preserves the betweenness relation and maps any branch $B$ of $D_{n_{1}}$ to a branch of $D_{n_{2}}$ whose color is the image via $\chi$ of the color of $B$. Let $F_{n}$ be an increasing exhausting sequence of finite subsets of $\mathrm{Br}(D_{n_{1}})$ such that $F_{0}$ contains just a point $b_{0}$ and $F_{n+1}\setminus F_{n}$ is a point in $[F_{n}]$ or a point whose projection is in $[F_{n}]$. The map $\tau$ is defined inductively as follows. Let $\tau(b_{0})$ be any branch point of $D_{n_{2}}$. Assume that $\tau$ has been defined on $F_{n}$, is a partial dendrite homomorphism (Definition 4.2) and, for any $b\neq b^{\prime}$ in $F_{n}$, is such that $k_{2}(c_{\tau(b)}(\tau(b^{\prime})))=\chi(k_{1}(c_{b}(b^{\prime})))$. Now let $b\in F_{n+1}\setminus F_{n}$ and let $p$ be its projection on $[F_{n}]$. If $p\in F_{n}$ then we choose a point $b_{2}$ in the branch $B$ of $D_{n_{2}}$ at $\tau(p)$ such that $k_{2}(B)=\chi(k_{1}(c_{p}(b)))$ and $k_{2}(c_{b_{2}}(\tau(p)))=\chi(k_{1}(c_{b}(p))$ and we define $\tau(b)=b_{2}$. If $p\notin F_{n}$ then $p=b\in[F_{n}]$ and there are $b_{1},b_{1}^{\prime}\in F_{n}$ such that $b\in[b_{1},b_{1}^{\prime}]$ and the interior of $[b_{1},b_{1}^{\prime}]$ contains no points of $F_{n}$. By the kaleidoscopic property, there is $b_{2}\in[\tau(b_{1}),\tau(b^{\prime}_{1})]$ such that $k_{2}(c_{b_{2}}(\tau(b_{1})))=\chi(k_{1}(c_{b}(b_{1})))$ and $k_{2}(c_{b_{2}}(\tau(b^{\prime}_{1})))=\chi(k_{1}(c_{b}(b^{\prime}_{1})))$. We set $\tau(b)=b_{2}$. This defines $\tau$ on $\mathrm{Br}(D_{n_{1}})$ with the required properties, which thus extends continuously to an embedding $\mathcal{X}\colon D_{n_{1}}\to D_{n_{2}}$ by [DMW19, Proposition 2.4].
@@
As soon as $n\leq m$, the group $\mathrm{Sym}([n])$ embeds into $\mathrm{Sym}([m])$ as the pointwise stabilizer of $[m]\setminus[n]$ points. With this embedding, the conditions of Lemma 4.14 are satisfied, so the first part of the statement is proved.
@@
Note that $m\geq 5$, so $\mathrm{Alt}(m)$ is simple. Then it is not hard to see that every faithful action of $\mathrm{Alt}(m)$ on a tree must have a fixed vertex $v$ on whose adjacent edges the action is the natural action of $\mathrm{Alt}(m)$ (see [Tar23, Claim 6.3] for details). Since $n<m$, the vertex $v$ of $T_{n,\infty}$ must be of infinite degree, so the local action around $v$ must induce an embedding of $\mathrm{Alt}(m)$ into $\mathrm{Aut}(S)$. This is impossible: if such an embedding existed, $\mathrm{Alt}(m)$ would either embed into $\mathrm{Aut}(O)$ or have an index-$2$ subgroup that embeds into $\mathrm{Aut}(O)$, both of which cannot happen because finite subgroups of $\mathrm{Aut}(O)$ must be cyclic. Thus, $\mathrm{Alt}(m)$ does not embed into $\mathcal{U}(\mathrm{Sym}([n]),\mathrm{Aut}(S))\simeq\mathrm{Homeo}(\mathcal{R}_{n})$, as needed. ∎
Corollary 1.30 and Proposition 4.15 immediately imply that, assuming that $\{n,m\}\neq\{3,4\}$, $\mathrm{Homeo}(\mathcal{R}_{n})\simeq\mathrm{Homeo}(\mathcal{R}_{m})$ if and only if $n=m$. For the case $n=3,m=4$, the strategy of Proposition 4.15 does not apply because $\mathrm{Alt}(4)$ is not simple. Indeed, [DM80] describes a faithful action of $\mathrm{Sym}([4])$ on a $3$-regular tree, so we cannot rule out that $\mathrm{Aut}(4)$ embeds into $\mathrm{Homeo}(\mathcal{R}_{3})$. Hence, the following questions remain open.
@@
Note that Proposition 4.15 does not say that there are no embeddings of $\mathrm{Homeo}(\mathcal{R}_{3})$ into $\mathrm{Homeo}(\mathcal{B})$. And indeed there are such embeddings:
@@
Embeddings between groups acting on the basilica and the airplane is not new behavior: as shown in [Tar24b, Section 10], the basilica rearrangement group $T_{B}$ embeds into the airplane rearrangement group $T_{A}$ and the overall idea is ultimately not dissimilar to that of Corollary 4.17.
@@ 4.5 Embeddings and non-embeddings among the lamination automorphism groups
Moreover (and perhaps more surprisingly), all of these groups embed into $\mathrm{Homeo}^{+}(\mathcal{B})$. Recalling again that $\mathrm{Homeo}^{+}(\mathcal{R}_{n})\simeq\mathcal{U}(\mathrm{Cyc}(n),\mathrm{Aut}(O))$ by Theorem 3.4, this can be seen as follows:
@@ 5.1 Quasi-isometry types of the groups
Thanks to Corollaries 2.18, LABEL: and 3.8, we know that both $\mathrm{Homeo}(\mathcal{A})$ and $\mathrm{Homeo}^{+}(\mathcal{A})$ are oligomorphic. In particular they are Roelcke precompact. See [Ros22, §3.1] for details. ∎
@@
Now let us prove that $\mathrm{Homeo}(\mathcal{R}_{n})$ and $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ are quasi-isometric to $T$. Since these groups act transitively on the set of (unoriented) edges of $T$, we can rely on the argument given in [Ros22, Example 6.35]. The fact that $T$ is quasi-isometric to $T_{\infty}$ follows from the following lemma. ∎
@@
For $n=0$, we set $f_{0}(v_{0})=u_{0}$. We extend $f_{0}$ to $f_{1}$ by choosing any bijection between the infinite countable sets $V_{1}\setminus\{v_{0}\}$ and $U_{1}\setminus\{u_{0}\}$. Assume that $f_{n}$ has been defined for some $n\geq 1$. We define $f_{n+1}$ to be $f_{n}$ on $V_{n}$. If $n$ is odd, for each vertex $v$ in $V_{n}\setminus V_{n-1}$, we choose a surjective map between neighbors of $v$ not in $V_{n}$ and the $k-1$ neighbors of $f(v_{n})$ not in $U_{n}$. This defines $f_{n+1}$ on the 1-neighborhood of $v$ and we do the same for all other vertices in $V_{n}\setminus V_{n-1}$.
@@ 5.2 Property (T) and the Haagerup property
Recall that a topological group $G$ has strong Property (T) if $G$ has a Kazhdan pair $(F,\varepsilon)$ with $F$ finite (see [BdlHV08] for details about Property (T)). In the opposite direction, the Haagerup property has been extended from locally compact groups to general topological group $G$ in [Ros17]. A topological group $G$ has the Haagerup property if $G$ has a continuous action by isometries on a Hilbert space $\mathcal{H}$ such that there is an orbit map that is a coarse embedding from $G$ to $\mathcal{H}$ (i.e., the preimages of bounded subsets are coarsely bounded). These properties are opposite in the sense that a topological group with both property (T) and the Haagerup property is coarsely bounded.
@@
Thanks to the GNS construction, the Haagerup property is equivalent to the existence of a continuous coarsely proper function conditionally of negative type. See [BdlHV08, §2.10] for explanations.
@@
For the airplane, thanks to the identification of both $\mathrm{Homeo}(\mathcal{A})$ and $\mathrm{Homeo}^{+}(\mathcal{A})$ with oligomorphic kaleidoscopic groups (Corollaries 2.18, LABEL: and 3.8), strong property (T) follows immediately from [ET16, Theorem 1.1].
Now consider the rabbits. Let $d$ be the distance function on $T_{\infty}$ or $T_{n,\infty}$ and $v$ some fixed vertex in this tree, the map $f\colon g\mapsto d(g(v),v)$ is a continuous, metrically proper and conditionally of negative type function [BdlHV08, C.2.2.(iii)] and thus gives rise to a continuous proper affine isometric on some Hilbert space (that is the Hilbert space of square-integrable functions on the countable set of oriented edges). ∎
By Theorem 5.5, the group $\mathrm{Homeo}^{+}(\mathcal{A})$ is a new example of an uncountable group with property (T) acting non-elementarily on the circle, such as that studied in [Duc23a].

### v2

@@ 4.2 Embedding trees into dendrites
For every $n\geq 3$, there exist color-preserving topological embeddings $\Phi\colon\overline{T_{n}}\to D_{n}$ such that the closure of $\bigcup_{\{v,w\}\in E(T_{n})}\overline{C_{\Phi(v),\Phi(w)}}$ is $D_{n}$.
@@
Let $\Phi_{*}\colon V(T_{n})\to\mathrm{Br}(D_{n})$ be defined as $\Phi_{m}(v)$ for $m$ such that $v\in B_{v_{0}}(m)$. For each edge $\{v,w\}$ of $T_{n}$, extend $\Phi_{*}$ on the corresponding arc $[v,w]$ of $\overline{T_{n}}$ to be any homeomorphism from $[v,w]$ to the unique arc $\Phi_{*}(v),\Phi_{*}(w)]$ of $D_{n}$. Thanks to condition (1) on each $\Phi_{m}$, this defines an embedding $\overline{T_{n}}\setminus\partial(T_{n})\to D_{n}\setminus\partial D_{n}$, which thus extends to an embedding $\Phi\colon\overline{T_{n}}\to D_{n}$. Thanks to condition (2) on the maps $\Phi_{m}$, the embedding $\Phi$ is color-preserving in the sense of Definition 4.7.
@@ 4.3 Embedding universal groups into kaleidoscopic groups
Given $g\in\mathcal{U}(N)$, we want to build $\varphi(g)\in\mathcal{K}(N)$. We start by considering $f\coloneqq\Phi\circ g\circ\Phi^{-1}$ on $\Phi\left(V(T_{n})\cup\partial T_{n}\right)$. Note that $f$ is a partial dendrite homomorphism (Definition 4.2). We will now extend $f$ “canonically” everywhere else using Lemma 4.4 with $F=F^{\prime}=\Phi\left(V(T_{n})\cup\partial T_{n}\right)$, as explained below.
By hypothesis, the closure of $\bigcup_{\{v,w\}\in E(T_{n})}\overline{C_{\Phi(v),\Phi(w)}}$ is the entirety of $D_{n}$, so $\Phi$ must map end points of $\overline{T_{n}}$ to end points of $D_{n}$. Thus the set of accumulation points of $\Phi\left(V(T_{n})\cup\partial T_{n}\right)$ is precisely $\Phi(\partial T_{n})$ and so $x,y]\cap\Phi\left(V(T_{n})\cup\partial T_{n}\right)$ is finite for all cut points $x$ and $y$ of $D_{n}$. Moreover, a connected component of $D_{n}\setminus\Phi\left(V(T_{n})\cup\partial T_{n}\right)$ coincides with a connected component of $D_{n}\setminus\Phi\left(V(T_{n})\right)$. Finally, the set $\Phi\left(V(T_{n})\cup\partial T_{n}\right)$ is closed and center-closed, so it satisfies all the hypotheses of Lemma 4.4.
Consider the set $\Omega_{\Phi(V(T_{n}))}$ of connected components of $D_{n}\setminus\Phi(V(T_{n}))$, which we will denote by $\Omega$ for the sake of brevity. Recall that the boundary of each element of $\Omega$ in $D_{n}$ consists of at most two elements of $\Phi(V(T_{n}))$ by Remark 4.3. Actually, since $\Phi$ is such that the closure of $\bigcup_{\{v,w\}\in E(T_{n})}\overline{C_{\Phi(v),\Phi(w)}}$ is $D_{n}$, the boundary of each element $A$ of $\Omega$ must have two points which belong to $\Phi(V(T_{n}))$, otherwise $A$ would be left out of such set. For the sake of brevity, given an edge $\{v,w\}$ of $T_{n}$, let us write $C_{v,w}$ for $C_{\Phi(v),\Phi(w)}$, which is the element of $\Omega$ whose boundary is $\{\Phi(v),\Phi(w)\}$. Let $D_{v,w}$ denote the closure of $C_{v,w}$, which is $C_{v,w}\cup\{\Phi(v),\Phi(w)\}$. Since $D_{v,w}$ is homeomorphic to $D_{n}$, for all $\{v,w\}\in E(T_{n})$ we can apply Proposition 4.6 to find color-preserving homeomorphisms $\psi_{(v,w)}\colon D_{v,w}\to D_{n}$ such that
@@
Since $\Phi$ is color-preserving (Definition 4.7) and the restriction $c_{b_{0}}|_{\Phi(HT(T_{n}))}$ is a bijection $\Phi(HT(T_{n}))\to\widehat{b_{0}}$, we have
$k_{T}=k_{D}\circ c_{b_{1}}\circ\Phi\circ\tau,$
@@
It is likely that statements analogous to Lemmas 4.8, LABEL:, 4.9, LABEL: and 4.11 also hold for universal groups $\mathcal{U}(N,M)$ on biregular trees $T_{n,m}$ and generalized kaleidoscopic groups $\mathcal{K}(N,M)$ on Ważewski dendrites $D_{n,m}$. However, the groups $\mathcal{K}(N,M)$ have not been formally defined in the literature. They can be built by using two disjoint sets of colors $[n]$ and $[m]$ and fixing two local groups $N$ and $M$ acting on $[n]$ and $[m]$, respectively, as noted in [DMW19, Remark 3.10].
@@ 4.4 Embeddings and non-embeddings among the homeomorphism groups
In the case of dendrites and kaleidoscopic groups, let us see that again $\chi$ induces color-preserving embeddings $\mathcal{X}\colon D_{n_{1}}\hookrightarrow D_{n_{2}}$. We will first produce an injective map $\tau\colon\mathrm{Br}(D_{n_{1}})\to\mathrm{Br}(D_{n_{2}})$ that preserves the betweenness relation and maps any branch $B$ of $D_{n_{1}}$ to a branch of $D_{n_{2}}$ whose color is the image via $\chi$ of the color of $B$. Let $F_{n}$ be an increasing exhausting sequence of finite subsets of $\mathrm{Br}(D_{n_{1}})$ such that $F_{0}$ contains just a point $b_{0}$ and $F_{n+1}\setminus F_{n}$ is a point in $F_{n}]$ or a point whose projection is in $F_{n}]$. The map $\tau$ is defined inductively as follows. Let $\tau(b_{0})$ be any branch point of $D_{n_{2}}$. Assume that $\tau$ has been defined on $F_{n}$, is a partial dendrite homomorphism (Definition 4.2) and, for any $b\neq b^{\prime}$ in $F_{n}$, is such that $k_{2}(c_{\tau(b)}(\tau(b^{\prime})))=\chi(k_{1}(c_{b}(b^{\prime})))$. Now let $b\in F_{n+1}\setminus F_{n}$ and let $p$ be its projection on $F_{n}]$. If $p\in F_{n}$ then we choose a point $b_{2}$ in the branch $B$ of $D_{n_{2}}$ at $\tau(p)$ such that $k_{2}(B)=\chi(k_{1}(c_{p}(b)))$ and $k_{2}(c_{b_{2}}(\tau(p)))=\chi(k_{1}(c_{b}(p))$ and we define $\tau(b)=b_{2}$. If $p\notin F_{n}$ then $p=b\in[F_{n}]$ and there are $b_{1},b_{1}^{\prime}\in F_{n}$ such that $b\in[b_{1},b_{1}^{\prime}]$ and the interior of $b_{1},b_{1}^{\prime}]$ contains no points of $F_{n}$. By the kaleidoscopic property, there is $b_{2}\in[\tau(b_{1}),\tau(b^{\prime}_{1})]$ such that $k_{2}(c_{b_{2}}(\tau(b_{1})))=\chi(k_{1}(c_{b}(b_{1})))$ and $k_{2}(c_{b_{2}}(\tau(b^{\prime}_{1})))=\chi(k_{1}(c_{b}(b^{\prime}_{1})))$. We set $\tau(b)=b_{2}$. This defines $\tau$ on $\mathrm{Br}(D_{n_{1}})$ with the required properties, which thus extends continuously to an embedding $\mathcal{X}\colon D_{n_{1}}\to D_{n_{2}}$ by [DMW19, Proposition 2.4].
@@
As soon as $n\leq m$, the group $\mathrm{Sym}([n])$ embeds into $\mathrm{Sym}([m])$ as the pointwise stabilizer of $m]\setminus[n]$ points. With this embedding, the conditions of Lemma 4.14 are satisfied, so the first part of the statement is proved.
@@
Note that $m\geq 5$, so $\mathrm{Alt}(m)$ is simple. Then it is not hard to see that every faithful action of $\mathrm{Alt}(m)$ on a tree must have a fixed vertex $v$ on whose adjacent edges the action is the natural action of $\mathrm{Alt}(m)$ (see [Tar25, Claim 6.3] for details). Since $n<m$, the vertex $v$ of $T_{n,\infty}$ must be of infinite degree, so the local action around $v$ must induce an embedding of $\mathrm{Alt}(m)$ into $\mathrm{Aut}(S)$. This is impossible: if such an embedding existed, $\mathrm{Alt}(m)$ would either embed into $\mathrm{Aut}(O)$ or have an index-$2$ subgroup that embeds into $\mathrm{Aut}(O)$, both of which cannot happen because finite subgroups of $\mathrm{Aut}(O)$ must be cyclic. Thus, $\mathrm{Alt}(m)$ does not embed into $\mathcal{U}(\mathrm{Sym}([n]),\mathrm{Aut}(S))\simeq\mathrm{Homeo}(\mathcal{R}_{n})$, as needed. ∎
Corollary 1.30 and Proposition 4.15 imply that, assuming that $\{n,m\}\neq\{3,4\}$, $\mathrm{Homeo}(\mathcal{R}_{n})\simeq\mathrm{Homeo}(\mathcal{R}_{m})$ if and only if $n=m$. For the case $n=3,m=4$, the strategy of Proposition 4.15 does not apply because $\mathrm{Alt}(4)$ is not simple. Indeed, [DM80] describes a faithful action of $\mathrm{Sym}([4])$ on a $3$-regular tree, so we cannot rule out that $\mathrm{Aut}(4)$ embeds into $\mathrm{Homeo}(\mathcal{R}_{3})$. Hence, the following questions remain open.
@@
Note that Proposition 4.15 does not state that there do not exist embeddings of $\mathrm{Homeo}(\mathcal{R}_{3})$ into $\mathrm{Homeo}(\mathcal{B})$. And indeed there are such embeddings:
@@
Embeddings between groups acting on the basilica and the airplane is not new behavior: as shown in [Tar24a, Section 10], the basilica rearrangement group $T_{B}$ embeds into the airplane rearrangement group $T_{A}$ and the overall idea is ultimately not dissimilar to that of Corollary 4.17.
@@ 4.5 Embeddings and non-embeddings among the lamination automorphism groups
Moreover (and perhaps more surprisingly), these groups embed into $\mathrm{Homeo}^{+}(\mathcal{B})$. Recalling again that $\mathrm{Homeo}^{+}(\mathcal{R}_{n})\simeq\mathcal{U}(\mathrm{Cyc}(n),\mathrm{Aut}(O))$ by Theorem 3.4, this can be seen as follows:
@@ 5.1 Quasi-isometry types of the groups
Thanks to Corollaries 2.18, LABEL: and 3.8, both $\mathrm{Homeo}(\mathcal{A})$ and $\mathrm{Homeo}^{+}(\mathcal{A})$ are oligomorphic. In particular they are Roelcke precompact. See [Ros22, §3.1] for details. ∎
@@
Now let us prove that $\mathrm{Homeo}(\mathcal{R}_{n})$ and $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ are quasi-isometric to $T_{\infty}$. Since these groups act transitively on the set of (unoriented) edges of $T$, we can rely on the argument given in [Ros22, Example 6.34]. The fact that $T$ is quasi-isometric to $T_{\infty}$ follows from the following lemma. ∎
@@
For $n=0$, we set $f_{0}(v_{0})=u_{0}$. We extend $f_{0}$ to $f_{1}$ by choosing any bijection between the infinite countable sets $V_{1}\setminus\{v_{0}\}$ and $U_{1}\setminus\{u_{0}\}$. Assume that $f_{n}$ has been defined for some $n\geq 1$. We define $f_{n+1}$ to be $f_{n}$ on $V_{n}$. If $n$ is odd, for each vertex $v$ in $V_{n}\setminus V_{n-1}$, we choose a surjective map between neighbors of $v$ not in $V_{n}$ and the k1k-1 neighbors of $f(v_{n})$ not in $U_{n}$. This defines $f_{n+1}$ on the 1-neighborhood of $v$ and we do the same for all other vertices in $V_{n}\setminus V_{n-1}$.
@@ 5.2 Property (T) and the Haagerup property
Recall that a topological group $G$ has strong Property (T) if $G$ has a Kazhdan pair $(F,\varepsilon)$ with $F$ finite (see [BHV08] for details about Property (T)). In the opposite direction, the Haagerup property has been extended from locally compact groups to general topological group $G$ in [Ros17]. A topological group $G$ has the Haagerup property if $G$ has a continuous action by isometries on a Hilbert space $\mathcal{H}$ such that there is an orbit map that is a coarse embedding from $G$ to $\mathcal{H}$ (i.e., the preimages of bounded subsets are coarsely bounded). These properties are opposite in the sense that a topological group with both property (T) and the Haagerup property is coarsely bounded.
@@
Thanks to the GNS construction, the Haagerup property is equivalent to the existence of a continuous coarsely proper function conditionally of negative type. See [BHV08, §2.10] for explanations.
@@
For the airplane, thanks to the identification of $\mathrm{Homeo}(\mathcal{A})$ and $\mathrm{Homeo}^{+}(\mathcal{A})$ with oligomorphic kaleidoscopic groups (Corollaries 2.18, LABEL: and 3.8), strong property (T) follows immediately from [ET16, Theorem 1.1].
Now consider the rabbits. Let $d$ be the distance function on $T_{\infty}$ or $T_{n,\infty}$ and $v$ some fixed vertex in this tree, the map $f\colon g\mapsto d(g(v),v)$ is a continuous, metrically proper and conditionally of negative type function [BHV08, C.2.2.(iii)] and thus gives rise to a continuous proper affine isometric on some Hilbert space (that is the Hilbert space of square-integrable functions on the countable set of oriented edges). ∎
By Theorem 5.5, the group $\mathrm{Homeo}^{+}(\mathcal{A})$ is a new example of an uncountable group with property (T) acting non-elementarily on the circle, such as that studied in [Duc23].